OpenStudy (anonymous):

Find the solution to the equation 64^(3 + x) = 4^2x.

OpenStudy (anonymous):

64^(3 + x) = 4^2x.

OpenStudy (anonymous):

x= -9

OpenStudy (ash2326):

We have \[64^{(3+x)}=4^{2x}\] We know \(64=4^3\) so \[(4^3)^{3+x}=4^{2x}\] We have now \[4^{9+3x}=4^{2x}\] Equating the powers of 4 both sides, we get \[9+3x=2x\] subtract both sides by 9 \[3x=2x-9\] subtract both sides by 2x, we get \[x=-9\]